5. Problem 9.07 PREFERRED STOCK RATE OF RETURN What will be the nominal rate of return on a perpetual preferred stock with a $100 par value, a stated dividend of 10% of par, and a current market price of (a) $54.00, (b) $89.00, (c) $104.00, and (d) $133.00? Round your answers to two decimal places. b. C. d.
Answer #1
Dividend = 10% of par
Dividend = 10% of 100
Dividend =10
Using the dividend discount model
a dividend that has a zero growth rate and is like a perpetuity

Price of preffered stock = Dividend/R

where R is the nominal rate of return.

a) when price is $54
54 = 10/R
R = 10/54
R = .1852 or 18.52%
b) when price is $89
89 = 10/R
R = 10/89
R = .1124 or 11.24%
c) when price is $104
104 = 10/R
R = 10/104
R = .09615 or 9.61%
d) when price is $133
133 = 10/R
R = 10/133
R = .0752 or 7.52%